OpenStudy (anonymous):

How do you solve for x in the following equation? 3x^2+36x+120=12

OpenStudy (anonymous):

3x²+36x+120=12 See that every term is dividable by 3?: x²+12x+40=4 Subtracting 4 from both sides: x²+12x+36=0 Now we need to find two numbers that multiplied become 26 and when added become 12: (x+6)(x+6)=0 Divide both side by (x+6): x+6=0 x=-6.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You can always check to see if the answer is correct also by plugging in. x²+12x+40=4 (-6)^2-72+40=4 36-72+40=4 36-22= 4 4 = 4 So -6 is correct.